baseball diamond is 90 feet to 1st and 90 feet to 2nd. Aprroximate direct distance from home plate to second base.
Answer by edjones(8007)

A baseball diamond is a square.
The distance from home to 2nd is the hypotenuse of an equilateral right triangle.
90^2+90^2=c^2
8100+8100=16200
sqrt(16200)=127.3'
.
Ed